a mass of 50 kg with an initial velocity of 100m/s is thrown vertically upward after 5s by explosion it is divided into two parts one part of mass 20 kg flies off vertically upward with velocity of 150 m/s what is the velocity of the other part 

Grade:11

1 Answers

Moazzam Ahmad
6 years ago
In the above Question There is application of law of conservation of momentum,
M = 50 kg which break into m1 = 20 kg and m2 = 30 kg. , Initial Velocity of M = Vi = 100 ms-1 and Velocity of m1 = v1 =150 ms-1
Velocity of m2 =  v2 = ?    
First Find Velocity after 5 sec as velocity decreases due to gravitational pull.
Apply Newtons First law of motion.
V = Vi – gt ( take upward motion as +ve so acc. due to gravity will be –ve)
V = 100 – (10 x 5)
V = 50 ms-1
According to law of conservation of momentum.
M x V = (m1 x v1) + (m2 x v2)
50 x 50 = (20 x 150) +(30 x v2)
2500 = 3000 +(30 x v2)
-50/3 ms-1 = v2.
v2 = -50/3 ms-1 (-ve means its direction is opposite to the direction of the 20 kg mass)